OpenStudy (anonymous):

2log(x-3)+1 = 5 Could you break it down step by step? I'm a bit confused on how to do this algebraically

OpenStudy (callisto):

To solve x?

OpenStudy (anonymous):

yes, i presume

OpenStudy (callisto):

Okay, first, subtract both sides by 1, what do you get?

OpenStudy (anonymous):

you get 2log(x-3) = 4

OpenStudy (callisto):

Right, now divide both sides by 2, what do you get?

OpenStudy (anonymous):

log(x-3) = 2

OpenStudy (anonymous):

haha, that's fine :)

OpenStudy (calculusfunctions):

Don't mind me. @Callisto seems to be doing a fine job teaching so I'll just observe if you don't mind? I'll only respond if you ask.

OpenStudy (callisto):

Yup! Now, take anti-log for both sides. What do you get? Note: anti-log = inverse of log For example, logx = 1 => 10^(logx) = 10^1 => x = 10^1 = 10

OpenStudy (anonymous):

yikes, haha, i'm not for sure. I briefly remember anti-logs but I'm not for sure what do with them

OpenStudy (callisto):

Do you understand that example?

OpenStudy (anonymous):

i'm starting to get it. logx = 1. I understand that but not after that

OpenStudy (calculusfunctions):

If I may, @Callisto perhaps he'd @Egirl01 would understand if you asked her to change the logarithm form to exponential form. @Egirl01 do you know how to do that.

OpenStudy (anonymous):

No, I'm sorry

OpenStudy (anonymous):

give me one second

OpenStudy (anonymous):

\(\large b^x=y \rightarrow log_by=x \)

OpenStudy (anonymous):

log^2 = x-3?

OpenStudy (callisto):

logx = 1 is actually just an example... ---------------------------------- Intuitively, log a number is the value of the power of the base Let say, I have a number 100, which can be rewritten as 10^2 Now, I take log (base 10) of that number (100), so I'll get 2, since 10 to the power of 2 give me 100. (Sorry if I make it even worse...)

OpenStudy (anonymous):

No, I understand the piece. so if its simply log, then its naturally log10, and the anti-log of log10 is 1/10 right?

OpenStudy (callisto):

anti-log of log 10 is 10. \[\log^{-1}(log10) = 10\] anti-log is the inverse of log.

OpenStudy (anonymous):

Alright. So then it'd be 10 (x-3) = 2?

OpenStudy (callisto):

No.... How did you get that?

OpenStudy (anonymous):

the anti-log of log 10 is 10, so log(x-3) = 2 would turn to 10(x-3) = 2? Obviously that is wrong. I'm a bit lost but you can continue onto the next part of solving the equation. I'll eventually figure what I'm doing wrong

OpenStudy (callisto):

\[log(x-3) = 2\]Take anti-log for both sides: \[\log^{-1}(log(x-3)) = \log^{-1}(2)\] Perhaps you can simplify the left first.

OpenStudy (anonymous):

I still don't quite follow, but continue on. I'll understand eventually :)

OpenStudy (callisto):

Nope.. It's your turn to work on it... Which part you don't understand ?

OpenStudy (anonymous):

how exactly I create the anti-log from log(x-3)

OpenStudy (callisto):

Because you are doing something on both sides... You create anti-log to undo the work of log...

OpenStudy (anonymous):

So the anti-log of log10 is 10, right? and the anti-log of 2 is what?

OpenStudy (callisto):

Before answering you question, could you please answer few questions first?

OpenStudy (anonymous):

okay

OpenStudy (callisto):

What is log10?

OpenStudy (anonymous):

log10 =1

OpenStudy (callisto):

What is log100?

OpenStudy (anonymous):

log100 = 2

OpenStudy (callisto):

Can we tell me what number do I have to take log on in order to get 1?

OpenStudy (anonymous):

10?

OpenStudy (callisto):

Yes. So, what number do I have to take log on in order to get 2?

OpenStudy (anonymous):

100

OpenStudy (callisto):

Yes! How did you get this answer?

OpenStudy (anonymous):

because you inversed it?

OpenStudy (callisto):

Yes! So, now, can you answer your question? what is \(\log^{-1}2\)?

OpenStudy (anonymous):

It is 100

OpenStudy (callisto):

Yes!

OpenStudy (anonymous):

So 10(x-3) = 100?

OpenStudy (callisto):

No. Another set question questions... What is log10 (again)?

OpenStudy (anonymous):

log 10 is 1

OpenStudy (callisto):

You got it right for the right side, it's wrong for the left.

OpenStudy (callisto):

What is \(\log^{-1}log(10)\)?

OpenStudy (anonymous):

10

OpenStudy (callisto):

Yes. What is \(\log^{-1}(log100)\)?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

100?

OpenStudy (callisto):

Yes!

OpenStudy (callisto):

What is \(\log^{-1} (log1000)\)?

OpenStudy (anonymous):

1000

OpenStudy (callisto):

So, we can see that \(\log^{-1}\) of log a number is the number itself, agree?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

So, \(\log^{-1} (\log y) = y\), agree?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

So, what is \(\log^{-1} (log(x-3))\)?

OpenStudy (anonymous):

x-3?

OpenStudy (callisto):

Yes!

OpenStudy (callisto):

Can you simplify the following now? \[\log^{-1}(log(x-3)) = \log^{-1}(2)\]

OpenStudy (anonymous):

x-3 = 100?

OpenStudy (callisto):

Yes! Now can you solve it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x = 103?

OpenStudy (callisto):

Nice :D and Yes! Do you understand how to solve this type of question now?

OpenStudy (anonymous):

yes! Thank you soooo much for your help! I really appreciate it!

OpenStudy (callisto):

You're welcome :)